# BatteryML: An Open-source Platform for Machine Learning on Battery Degradation

### v4

@@ B.3 State of Charge Estimation
Contrasting with the results for RUL and SOH prediction, SOC benchmark results reveal that LightGBM Ke et al. (2017) consistently surpasses other methodologies in the majority of tasks, thereby positioning tree models as the prevailing state-of-the-art for SOC prediction. However, all methods exhibit suboptimal performance on the RWTH dataset, suggesting that current approaches still grapple with adapting to the full spectrum of battery aging patterns. Additionally, linear models persistently outperform deep learning models, implying that both the input features and the network architecture necessitate further optimization to fully exploit the potential of deep learning models in SOC prediction tasks.
Appendix C Ablation Study
@@
C.1 Feature space ablation
@@ C.2 Hyperparameter ablation
Figure 7 illustrates the impact of hidden dimensions on the performance of deep models. In Fig 7a, an increase in hidden dimensions results in higher variance without a significant change in mean prediction value, suggesting reduced robustness for larger dimensions in the MIX data source.
@@
Appendix D Flexible Extensions Based on BatteryML
D.1 Pipeline Interface for Efficient Replication
@@
        - name: ’LogScaleDataTransformation’
        - name: ’ZScoreDataTransformation’
@@
Almost all modules in BatteryML can be configured using the configuration file. Code D.1 demonstrates an example that specifies “Variance” model on MATR1 dataset. Here we introduce these components in detail.
@@
Module feature_transformation and label_transformation.   These two module is responsible for post-processing features and labels before training. Examples include normalization and data augmentation. In addition to common transformations, we also implemented a sequential transformation wrapper that allows flexible compositions of transformations. For example, in Code D.1 we transforms the label by first converting to log scale and then applying z-score normalization. The corresponding code path is batteryml/data/transformation.
@@
Combining these modules together, researchers can employ the ‘Pipeline‘ of BatteryML to conduct experiments efficiently, as demonstrated in D.1. Please refer to the documentation for more examples.
@@
D.2 Unified Data Format for Public and Custom Data
@@
BatteryML abstracts a general representation BatteryData for the diverse battery data records in existing studies. This abstraction lies at the core of BatteryML, which organizes the arbitrary form of battery data into a unified format, covering most public datasets and the output formats of typical battery cycling test equipment. Here we use MATR data source as an example for demonstration. As shown in Table 6, each BatteryData represents a battery cell. For each cell, meta information such as cathode material, anode material, etc., as well as charge-discharge cycle data and charge-discharge protocols are recorded. We assume that each cell is related to an array of charge-discharge cycles, which we store as a list of cycling records, as shown in Table  7. Each cycle comprises time-series records that include electrical signals such as voltage and current, and may also incorporate other parameters like temperature, resistance, mass, etc. Table  8 demonstrates the data entry method for charge-discharge protocols using a discharge protocol as an example. For instance, in the MATR dataset, the discharge is constant current, so the rate_in_C is fixed at 4C, the starting capacity is 1, the ending capacity is 0, and the cycle is continuous. BatteryML also provides a command line interface to conveniently organize existing public datasets into BatteryData, as shown in Code D.2. Currently BatteryML supports cell-level learning of batteries for degradation modeling or battery optimization. In the future, BatteryML will accommodate a wider range of real-world applications, including higher-level battery packages such as BMS records of electric vehicles and lower-level data such as half-cells and material properties. BatteryML will continually foster data exchange in the battery domain, reducing the barriers to data-driven battery modeling, thereby paving the way for future scientific research.
@@
D.3 Feature Interface for Domain Knowledge Incorporation
@@
To develop custom feature extractors, developers simply need to implement the process_cell function, which is designed to handle a specific cell. BatteryML efficiently manages the iteration through the batteries and organizes the data collation into structured datasets. Code D.3 demonstrates a practical example that calculates the Coulombic efficiency as input features.
@@
The FEATURE_EXTRACTORS registry manager will add the custom feature extractor into BatteryML. After we add this class into the import stack of BatteryML, coulombic efficiency feature can be used in configuration files for training, as shown in Code D.3.
@@
        - name: ’LogScaleDataTransformation’
        - name: ’ZScoreDataTransformation’
D.4 Preprocessing Interface for Custom Data Cleaning and Augmentation
@@
Code D.4 showcases an example of custom data preprocessor, where the input data is normalized by the minimum and maximum values. Code D.4 demonstrated an configuration example that uses this min-max-normalization feature for learning.
@@
        - name: ’LogScaleDataTransformation’
        - name: ’MinMaxDataTransformation’
D.5 Model Interface for Custom Battery Models
@@
Since lightgbm.LGBMRegressor already supports fit and predict for predictions, we can directly inherit the batteryml.model.SklearnModel class and pass the hyperparameters of the model through configuration file. Code D.5 demonstrates the example implementation for the RUL predictor using LightGBM. In this example, BatteryML will call the fit and predict method of LightgbmRULPredictor.model for the training and evaluation. Finally, the custom LightGBM model is registered to BatteryML after @MODELS.register() and imported during initialization.
@@
After defining the model, we can pass the hyperparameters of the RUL predictor through the configuration file, as shown in Code  D.5. The fields in the configuration file will be passed to the model as keyword arguments when initializing the predictor.
@@
        - name: ’LogScaleDataTransformation’
        - name: ’ZScoreDataTransformation’
Linear Regression.
Elastic Net.
Gaussian Process.
Principal Component Regression.
Partial Least Squares Regression.
Ridge Regression.
Support Vector Regression.
XGBoost Regression.
XGBoostRegressor is an implementation of the gradient boosting decision trees algorithm, designed for speed and performance, that is particularly suited for regression problems, providing a robust and versatile model capable of handling a variety of data types, achieving excellent predictive performance. We use their scikit-learn interface for implementation
Random Forest.
Dummy Regression.
This model uses the mean of the training samples as predictions, usually acting as a trivial baseline to indicate the error upper bound.
Multilayer Perceptron.
Long-Short-Term Memory.
Convolutional Neural Networks.
Transformer.

### v5

@@ C.3 State of Charge Estimation
Contrasting with the results for RUL and SOH prediction, SOC benchmark results reveal that LightGBM consistently surpasses other methodologies in the majority of tasks, thereby positioning tree models as the prevailing state-of-the-art for SOC prediction. However, all methods exhibit suboptimal performance on the RWTH dataset, suggesting that current approaches still grapple with adapting to the full spectrum of battery aging patterns. Additionally, linear models persistently outperform deep learning models, implying that both the input features and the network architecture necessitate further optimization to fully exploit the potential of deep learning models in SOC prediction tasks.
Appendix D Ablation Study
@@
D.1 Feature space ablation
@@ D.2 Hyperparameter ablation
Figure 7 illustrates the impact of hidden dimensions on the performance of deep models. In Figure 7a, an increase in hidden dimensions results in higher variance without a significant change in mean prediction value, suggesting reduced robustness for larger dimensions in the MIX data source.
@@
Appendix E Flexible Extensions Based on BatteryML
E.1 Pipeline Interface for Efficient Replication
@@
- name: ’LogScaleDataTransformation’
- name: ’ZScoreDataTransformation’
@@
Almost all modules in BatteryML can be configured using the configuration file. Code E.1 demonstrates an example that specifies “Variance” model on MATR1 dataset. Here we introduce these components in detail.
@@
Module feature_transformation and label_transformation.   These two module is responsible for post-processing features and labels before training. Examples include normalization and data augmentation. In addition to common transformations, we also implemented a sequential transformation wrapper that allows flexible compositions of transformations. For example, in Code E.1 we transforms the label by first converting to log scale and then applying z-score normalization. The corresponding code path is batteryml/data/transformation.
@@
Combining these modules together, researchers can employ the ‘Pipeline‘ of BatteryML to conduct experiments efficiently, as demonstrated in Code E.1. Please refer to the documentation for more examples.
@@
E.2 Unified Data Format for Public and Custom Data
@@
BatteryML abstracts a general representation BatteryData for the diverse battery data records in existing studies. This abstraction lies at the core of BatteryML, which organizes the arbitrary form of battery data into a unified format, covering most public datasets and the output formats of typical battery cycling test equipment. Here we use MATR data source as an example for demonstration. As shown in Table 6, each BatteryData represents a battery cell. For each cell, meta information such as cathode material, anode material, etc., as well as charge-discharge cycle data and charge-discharge protocols are recorded. We assume that each cell is related to an array of charge-discharge cycles, which we store as a list of cycling records, as shown in Table 7. Each cycle comprises time-series records that include electrical signals such as voltage and current, and may also incorporate other parameters like temperature, resistance, mass, etc. Table 8 demonstrates the data entry method for charge-discharge protocols using a discharge protocol as an example. For instance, in the MATR dataset, the discharge is constant current, so the rate_in_C is fixed at 4C, the starting capacity is 1, the ending capacity is 0, and the cycle is continuous. BatteryML also provides a command line interface to conveniently organize existing public datasets into BatteryData, as shown in Code E.2. Currently BatteryML supports cell-level learning of batteries for degradation modeling or battery optimization. In the future, BatteryML will accommodate a wider range of real-world applications, including higher-level battery packages such as BMS records of electric vehicles and lower-level data such as half-cells and material properties. BatteryML will continually foster data exchange in the battery domain, reducing the barriers to data-driven battery modeling, thereby paving the way for future scientific research.
@@
E.3 Feature Interface for Domain Knowledge Incorporation
@@
To develop custom feature extractors, developers simply need to implement the process_cell function, which is designed to handle a specific cell. BatteryML efficiently manages the iteration through the batteries and organizes the data collation into structured datasets. Code E.3 demonstrates a practical example that calculates the Coulombic efficiency as input features.
@@
The FEATURE_EXTRACTORS registry manager will add the custom feature extractor into BatteryML. After we add this class into the import stack of BatteryML, coulombic efficiency feature can be used in configuration files for training, as shown in Code E.3.
@@
- name: ’LogScaleDataTransformation’
- name: ’ZScoreDataTransformation’
E.4 Preprocessing Interface for Custom Data Cleaning and Augmentation
@@
Code E.4 showcases an example of custom data preprocessor, where the input data is normalized by the minimum and maximum values. Code E.4 demonstrated an configuration example that uses this min-max-normalization feature for learning.
@@
- name: ’LogScaleDataTransformation’
- name: ’MinMaxDataTransformation’
E.5 Model Interface for Custom Battery Models
@@
Since lightgbm.LGBMRegressor already supports fit and predict for predictions, we can directly inherit the batteryml.model.SklearnModel class and pass the hyperparameters of the model through configuration file. Code E.5 demonstrates the example implementation for the RUL predictor using LightGBM. In this example, BatteryML will call the fit and predict method of LightgbmRULPredictor.model for the training and evaluation. Finally, the custom LightGBM model is registered to BatteryML after @MODELS.register() and imported during initialization.
@@
After defining the model, we can pass the hyperparameters of the RUL predictor through the configuration file, as shown in Code E.5. The fields in the configuration file will be passed to the model as keyword arguments when initializing the predictor.
@@
- name: ’LogScaleDataTransformation’
- name: ’ZScoreDataTransformation’